OpenStudy (anonymous):

What is the value of x in the equation below? 14x - 6 = -3 + 2x + 5

OpenStudy (australopithecus):

group like terms

OpenStudy (australopithecus):

then factor out x

OpenStudy (zepp):

14x - 6 = -3 + 2x + 5 +6 +6 14x = -3 + 2x + 5 + 6 14x = 2x + 8 14x - 2x = 8 12x = 8 x = 8/12